DETECTION OF THE NEAR-INFRARED SPECTRUM OF CH_{2}^{+}

\underline{JENNIFER L. GOTTFRIED} and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

The molecular ion CH_{2}^{+} is of special interest because of its quasilinearity (resulting from a small barrier to linearity, 1089 cm^{−1}) and the strong interaction of its ground and first excited states due to the Renner-Teller effect. At linearity the ground state is a doubly degenerate ^{2}Π_{u} state that splits into X̃^{2}A_{1} and Ã^{2}B_{1} states as the molecule bends. Although in the ground state the molecule is a near-prolate asymmetric rotor (^{}H̸CH=140^{∘}), the excited state equilibrium geometry is linear. Both the initial detection of the infrared spectrum of the ν_{3} band^{a}^{,}^{b} and a more recent PFI-ZEKE study^{c} confirm the bent nature of CH_{2}^{+} in the ground state. This work presents the first experimental spectroscopic characterization of the Ã^{2}B_{1} state.

Using a Ti:sapphire laser, we have observed four new absorption bands in the near-infrared region (11,000-13,000 cm^{−1}). The comparison between spectra recorded in He/CH_{4} and He/CH_{4}/H_{2} liquid-nitrogen cooled positive column discharges was used to identify these transitions, the strongest of which had a signal-to-noise ratio of ∼100. A least-squares fit of 57 transitions from the Ã(0,3,0)^{1}←X̃(0,0,0)^{0} band has given preliminary values for the excited state molecular constants B=7.140(22) and D_{N}=-0.00034(16). The fit of 46 lines from the Ã(0,4,0)^{0}←X̃(0,0,0)^{1} band resulted in the following excited state constants: B=6.796(56) and D_{N}=-0.00170(42). Recent ab initio predictions^{d} of the rovibronic spectra of CH_{2}^{+} show good agreement with the observed spectrum (△ν_{o−c}=-53 cm^{−1} for the Ã(0,3,0)^{1}←X̃(0,0,0)^{0} band and △ν_{o−c}=-18 cm^{−1} for the Ã(0,4,0)^{0}←X̃(0,0,0)^{1} band). Assignment of the Ã(0,3,0)^{2}←X̃(0,0,0)^{1} and Ã(0,3,0)^{3}←X̃(0,0,0)^{2} bands is underway.

aM. Rösslein, C.M. Gabrys, M.-F. Jagod, and T. Oka, J. Mol. Spectrosc. \underline{153}, 738 (1992)

bP. Jensen, M. Brumm, W.P. Kraemer, and P.R. Bunker, J. Mol. Spectrosc. \underline{172}, 194 (1995)

cS. Willitsch and F. Merkt, J. Chem. Phys. \underline{118}, 2235 (2003)

dP.R. Bunker, M.C. Chan, W.P. Kraemer, and P. Jensen, Chem. Phys. Lett. \underline{341}, 358 (2001)